OpenStudy (anonymous):

Simplify the square root of 72 a. square root of 12 b. 6* square root of 2 c. 2* square root of 6 d. 36* square root of 2

OpenStudy (anonymous):

\[\sqrt{72}=\sqrt{36}*\sqrt{2}=6\sqrt{2}\]

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

@Posideon : how do you do the square root symbol lol.

OpenStudy (anonymous):

use the equation thing at the bottom